please sir solve this question divide rs 21 in to two parts so that they are 3\4 of the other
Answer by skartikey(21)

let two parts are x and 21-x
we are given
(3/4)x=21-x
3x=84-4x
3x+4x=84
7x=84
x=12
21-x=21-12=9
two parts are 12 and 9